# F-Fidelity: A Robust Framework for Faithfulness Evaluation of Explainable AI

### v1

@@ Abstract
Recent research has developed a number of eXplainable AI (XAI) techniques, such as gradient-based approaches, input perturbation-base methods, and black-box explanation methods. Although extracting meaningful insights from deep learning models, how to properly evaluate these XAI methods remains an open problem. The most widely used approach is to perturb or even remove what the XAI method considers to be the most important features in an input and observe the changes in the output prediction. This approach although efficient suffers the Out-of-Distribution (OOD) problem as the perturbed samples may no longer follow the original data distribution. A recent method RemOve And Retrain (ROAR) solves the OOD issue by retraining the model with perturbed samples guided by explanations. However, the training may not always converge given the distribution difference. Furthermore, using the model retrained based on XAI methods to evaluate these explainers may cause information leakage and thus lead to unfair comparisons. We propose Fine-tuned Fidelity (F-Fidelity), a robust evaluation framework for XAI, which utilizes i) an explanation-agnostic fine-tuning strategy, thus mitigating the information leakage issue and ii) a random masking operation that ensures that the removal step does not generate an OOD input. We designed controlled experiments with state-of-the-art (SOTA) explainers and their degraded version to verify the correctness of our framework. We conducted experiments on multiple data structures, such as images, time series, and natural language. The results demonstrate that F-Fidelity significantly improves upon prior evaluation metrics in recovering the ground-truth ranking of the explainers. Furthermore, we show both theoretically and empirically that, given a faithful explainer, F-Fidelity metric can be used to compute the sparsity of influential input components, i.e., to extract the true explanation size.
@@ 1 Introduction
Despite extracting useful insights about model decisions, how to faithfully evaluate and compare explanation methods remains an open challenge. There have been some existing efforts to address this issue. The preliminary works create datasets with “ground-truth” explanations for XAI evaluation, such as the MS-CoCo dataset in CV, (Lin et al., 2014), the Mutag dataset in graph (Debnath et al., 1991), and the e-SNLI dataset in NLP (Camburu et al., 2018). However, these datasets are still limited to certain tasks and cannot be generalized. More importantly, these “ground-truth” explanations are created based on humans’ understanding of a data sample, which may not reflect an ML model’s decision-making processes, and thus may still give unfaithful evaluations. More recent works propose the removal strategy, which does not rely on “ground-truth” explanations (Zhou et al., 2016; Selvaraju et al., 2017; Rong et al., 2022; Hooker et al., 2019; Madsen et al., 2023; 2022; Hase et al., 2021; Zheng et al., 2023; Yuan et al., 2022). Technically speaking, the removal strategy removes certain parts of an input that are deemed as (non-)important and records the changes in the model prediction. A larger drop in accuracy when removing the important parts reflects the removed parts are indeed important and thus validates the faithfulness of the corresponding explanation method. Specific metrics designed based on the removal strategy include the Most Relevant First (MoRF) and Least Relevant First (LeRF) (Samek et al., 2016; Yuan et al., 2022) in CV and Importance Measures (Madsen et al., 2021; 2023) in NLP. However, these removal-based explanations suffer from the Out-Of-Distribution (OOD) issue (Hooker et al., 2019), as the perturbed samples with removed features may no longer follow the original distribution. As a result, the model predictions of the perturbed samples are unreliable, regardless of the explanation faithfulness.
@@
To address the OOD problem, RemOve And Retrain (ROAR) (Hooker et al., 2019), proposed retraining the model for evaluating the explainer, where the model is trained using explanation-guided data augmentations. However, retraining the model based on the explainer output leads to information leakage and convergence issues (Rong et al., 2022). An additional issue, which has been overlooked in prior works, is that the removal strategy is highly dependent on explanation size ( or sparsity). That is, the size of the explanation part (or non-explanation part) that is removed affects the evaluation output. In the absence of ground-truth explanations, it is not clear what fraction of the input elements should be removed as part of the explanation, which leads to inconsistent evaluations.
@@
To address the aforementioned challenges, we introduce Fine-tuned Fidelity (F-Fidelity), a novel framework to evalute the performance of explainers. F-Fidelity leverages i) an explanation-agnostic fine-tuning strategy to prevent information leakage and ii) a controlled random masking operation to overcome the OOD issues observed in the application of prior evaluation metrics. The fine-tuning process employs stochastic masking operations, such as randomly dropping pixels in images, tokens in language, or time steps in time series data, to generate augmented training samples. This augmented data is then used to fine-tune a surrogate model. Unlike previous approaches such as ROAR (Hooker et al., 2019), our strategy effectively mitigates the risk of information leakage and label bias by using explanation-agnostic stochastic masks, while also offering improved efficiency as it eliminates the need to retrain the model for each individual explainer. During the evaluation phase, F-Fidelity implements a removal strategy that generates stochastic masks conditioned on the explainer output, designed to be in-distribution with respect to the masks used in the fine-tuning step, which mitigates the OOD issues observed in the application of prior evaluation metrics.
@@ 2 Preliminaries
This section introduces the notation and key background concepts used throughout the paper.
@@ 2.3 Quantifying the Performance of Explainers
A key challenge in explainability research is to quantify and compare the performance of various explainers. The performance of an explainer can be formally quantified in terms the total variation distance $d_{TV}(P_{Y|\mathbf{X}},P_{Y|\mathbf{X}\odot\mathbf{M}})$ as a function of the average explanation size $\mathbb{E}_{\mathbf{X}}(\|\mathbf{M}\|_{1})$. However, in most problems of interest, the underlying statistics $P_{\mathbf{X},Y}$ is not available, and hence direct evaluation of the aforementioned total variation distance is not possible. As discussed in the introduction, some datasets are accompanied by ground-truth explanations, which enables the use of measures such as AUC and IoU for evaluating the quality explainers. However, the ground-truth explanations are available only for a limited collection of datasets, and even when ground-truth explanations are available, they may not accurately reflect the model’s internal decision-making processes. To address the aforementioned issues, a widely used set of metrics have been proposed in the literature which are based on the removal strategy. In CV, two removal orders have been considered (Samek et al., 2016; Yuan et al., 2022), MoRF, where the removal process begins with the most influential pixels, and LeRF, where it starts with the least influential pixels. In the graph domain, an alternative removal strategy is used, where first the size of explanation subgraphs is determined according to a sparsity parameter, and then edges are removed either from the explanation subgraph of the desired size, or the non-explanation subgraph which is its complement. In this paper, we use metrics which are based on the removal strategy used in the graph domain, called the Fidelity (Fid) metric in the literature Pope et al. (2019); Bajaj et al. (2021); Yuan et al. (2022). Formally, given the input and label pair $(\mathbf{x},y)$ and a binary mask $\mathbf{m}$, the $Fidelity$ metrics are defined as follows:
$Fid^{+}(\psi)$ $=\frac{1}{n}\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{T}}\mathds{1}(y_{i}=f(\mathbf{x}_{i}))-\mathds{1}(y_{i}=f(\mathbf{x}_{i}-\mathbf{x}_{i}\odot\mathbf{m}_{i}))$ (1)
$Fid^{-}(\psi)$ $=\frac{1}{n}\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{T}}\mathds{1}(y_{i}=f(\mathbf{x}_{i}))-\mathds{1}(y_{i}=f(\mathbf{x}_{i}\odot\mathbf{m}_{i})),$ (2)
@@
where $\mathcal{T}$ is the dataset used for evaluating the performance of the explainer, $n$ is the size of the dataset, $\mathds{1}(\cdot)$ denotes the indicator function, and $\mathbf{m}_{i}=\psi(\mathbf{x}_{i})$ is the explanation corresponding to $\mathbf{x}_{i}$ produced by the explainer $\psi(\cdot)$.
@@
In R-Fidelity (Zheng et al., 2023), which considered the evaluation of GNN explainers, it was argued that the Fidelity metric highly relies on the robustness of the underlying classifier to removal of potentially large sections of the input, e.g.,the removal of a large subgraph explanation for $Fid^{+}$ or its complement for $Fid^{-}$. That is, the classifier should be robust to out-of-distribution (OOD) inputs for Fidelity metric to align with those of the (theoretically justified) total-variation-based metric discussed in the previous sections. To mitigate these OOD issues in the evaluation of the Fidelity metric, the following Robust Fidelity metrics (RFid) was introduced:
$RFid^{+}(\psi,\alpha^{+},s)$ $=\frac{1}{n}\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{T}}\mathds{1}(y_{i}=f(\mathbf{x}_{i}))-P(y_{i}=f(\chi^{+}(\mathbf{x}_{i},\alpha^{+},s))$ (3)
$RFid^{-}(\psi,\alpha^{-},s)$ $=\frac{1}{n}\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{T}}\mathds{1}(y_{i}=f(\mathbf{x}_{i}))-P(y_{i}=f(\chi^{-}(\mathbf{x}_{i},\alpha^{-},s)),$ (4)
@@
where $\chi^{+}(\mathbf{x}_{i},\alpha^{+},s)$ is a sampling function which randomly, uniformly, and independently removes $\lfloor s\alpha^{+}\rfloor$ elements from the $s$ highest scoring elements of $\mathbf{x}_{i}$ based on the scores produced by $\phi(\mathbf{x}_{i})$, and $\chi^{-}(\mathbf{x}_{i},\alpha^{-})$ removes $\lceil s\alpha^{-}\rceil$ fraction of the lowest scoring $td-s$ elements. It can be noted that if $\alpha^{+}=\alpha^{-}=1$, then the RFid metric reduces to the Fid metric. On the other hand, as $\alpha^{+}$ and $\alpha^{-}$ are decreased, fewer input elements are removed, hence requiring lower OOD robustness to ensure the accuracy of the evaluation output. In this work, we adapt the RFid metric for general, non-graphical domains, and introduce improvements in several directions. First, we show through extensive empirical evaluation that a specific finetuning process applied to the to-be-explained classifier enable significantly more precises evaluation of explainers using the RFid metric. Second, we show both theoretically and empirically, that by appropriate choice of the $\alpha^{+}$ and $\alpha^{-}$ parameters, one can extract the size of the ground-truth explanation by evaluating the RFid metric.
3 Robust Fidelity via Model Fine-tuning and Bounded Removals
@@
A partial solution in the graph domain addresses this issue by removing only an $\alpha^{+}$ fraction of the explanation subgraph and $\alpha^{-}$ fraction of the non-explanation subgraphs (Zheng et al., 2023). However, we argue that two issues degrade the evaluation quality of the RFid metric. First, if the original explanation size $\mathbf{M}$ is large (small), then removing an $\alpha^{+}$ ($\alpha^{-}$) of the explanation (non-explanation) part of the input, this would still yield OOD inputs. Second, the classifier may lack robustness and produce unreliable outputs even when the input is only slightly perturbed.
To address the first issue, we modify the RFid metric so that the size of the removed part of the input is upper-bounded by a constant fraction of the entire input, rather than a fraction of the explanation, which can vary depending on the learning task. That is, for a fixed $\beta\in[0,1]$, and RFid parameters $\alpha^{+}_{orig},\alpha^{-}_{orig}\in[0,1]$, we set the upper-bounded RFid parameters as
$\alpha^{+}=\min(\alpha^{+}_{orig},\frac{\beta td}{s})\quad\text{and}\quad\alpha^{-}=\min(\alpha^{-}_{orig},\frac{\beta td}{s}),$ (5)
so that the sampling functions $\xi^{+}$ and $\xi^{-}$ do not remove more than at most $\beta td$ elements of the input.
@@
To address the second issue, we fine-tune the classifier prior to applying the explanation method. This finetuning process is performed by taking into consideration the aforementioned upper bound on the size of the removed input. To elaborate, we introduce a stochastic mask generator $P_{\beta}:(t,d)\mapsto\mathbf{M}_{\beta}$, which takes a the input dimensions $(t,d)$ as input and outputs a mask $\mathbf{M}_{\beta}\in\{0,1\}^{t\times d}$ of size $\beta td$, i.e. with $\beta{td}$ non-zero elements. For instance, in image classification, the mask generator is designed to select random image patches for removal. Formally, we define the fine-tuning loss function as:
$L=\mathbb{E}_{\mathbf{X},Y}\left[\mathcal{L}\left(f\left(\mathbf{X}-P_{\beta}\odot\mathbf{X}\right),\ Y\right)\right]$ (6)
@@
where $\mathcal{L}$ is the loss function used during training (e.g., cross-entropy). The pipeline of our evaluation method, which incorporates the fine-tuning and upper-bounded removal techniques described above is provided in Algorithm 1. We denote the resulting evaluation metrics, which utilize the finetuning process and the $RFid^{+}$ and $RFid^{-}$ metrics with sampling rates which are truncated based on $\beta$ ( equation 5), as $FFid^{+}(\psi,\alpha^{+}_{orig},\beta,s)$ and $FFid^{-}(\psi,\alpha^{-}_{orig},\beta,s)$, respectively.
4 Empirical Verification of Robustness

### v2

@@ Abstract
Recent research has developed a number of eXplainable AI (XAI) techniques, such as gradient-based approaches, input perturbation-base methods, and black-box explanation methods. While these XAI techniques can extract meaningful insights from deep learning models, how to properly evaluate them remains an open problem. The most widely used approach is to perturb or even remove what the XAI method considers to be the most important features in an input and observe the changes in the output prediction. This approach, although straightforward, suffers the Out-of-Distribution (OOD) problem as the perturbed samples may no longer follow the original data distribution. A recent method RemOve And Retrain (ROAR) solves the OOD issue by retraining the model with perturbed samples guided by explanations. However, using the model retrained based on XAI methods to evaluate these explainers may cause information leakage and thus lead to unfair comparisons. We propose Fine-tuned Fidelity (F-Fidelity), a robust evaluation framework for XAI, which utilizes i) an explanation-agnostic fine-tuning strategy, thus mitigating the information leakage issue, and ii) a random masking operation that ensures that the removal step does not generate an OOD input. We also design controlled experiments with state-of-the-art (SOTA) explainers and their degraded version to verify the correctness of our framework. We conduct experiments on multiple data modalities, such as images, time series, and natural language. The results demonstrate that F-Fidelity significantly improves upon prior evaluation metrics in recovering the ground-truth ranking of the explainers. Furthermore, we show both theoretically and empirically that, given a faithful explainer, F-Fidelity metric can be used to compute the sparsity of influential input components, i.e., to extract the true explanation size. The source code is available at https://trustai4s-lab.github.io/ffidelity.
footnotetext: Corresponding authors
@@ 1 Introduction
Despite extracting useful insights about model decisions, how to faithfully evaluate and compare explanation methods remains an open challenge. There have been some existing efforts to address this issue. The preliminary works create datasets with “ground-truth” explanations for XAI evaluation, such as the MS-CoCo dataset in CV, (Lin et al., 2014), the Mutag dataset in graph (Debnath et al., 1991), and the e-SNLI dataset in NLP (Camburu et al., 2018). However, these datasets are still limited to certain tasks and cannot be generalized. More importantly, these “ground-truth” explanations are created based on humans’ understanding of a data sample, which may not reflect an ML model’s decision-making processes, and thus may still give unfaithful evaluations. More recent works propose the removal strategy, which does not rely on “ground-truth” explanations (Zhou et al., 2016; Selvaraju et al., 2017; Rong et al., 2022; Hooker et al., 2019; Madsen et al., 2023; 2022; Hase et al., 2021; Zheng et al., 2024; Yuan et al., 2022). Technically speaking, the removal strategy removes certain parts of an input that are deemed as (non-)important and records the changes in the model prediction. A larger drop in accuracy when removing the important parts reflects the removed parts as indeed important and thus validates the faithfulness of the corresponding explanation method. For example, in computer vision, removal means setting important pixels to zero/black; in natural language processing, it means replacing chosen tokens with “[MASK]” token or zero embeddings. Specific metrics designed based on the removal strategy include the Most Relevant First (MoRF) and Least Relevant First (LeRF) (Samek et al., 2016; Yuan et al., 2022) in CV and Importance Measures (Madsen et al., 2021; 2023) in NLP. However, these removal-based explanations suffer from the Out-Of-Distribution (OOD) issue (Hooker et al., 2019), as the perturbed samples with removed features may no longer follow the original distribution. As a result, the model predictions of the perturbed samples are unreliable, regardless of the explanation faithfulness.
@@
To address the OOD problem, RemOve And Retrain (ROAR) (Hooker et al., 2019), proposed retraining the model for evaluating the explainer, where the model is trained using explanation-guided removal. However, retraining the model based on the explainer output leads to an information leakage issue (Rong et al., 2022). An additional issue, which has been overlooked in prior works, is that the removal strategy is highly dependent on explanation size ( or sparsity). That is, the size of the explanation part that is removed affects the evaluation output. In the absence of ground-truth explanations, it is not clear what fraction of the input elements should be removed as part of the explanation, which leads to inconsistent evaluations.
@@
To address the aforementioned challenges, we introduce Fine-tuned Fidelity (F-Fidelity), a novel framework to evaluate the performance of explainers. F-Fidelity leverages i) an explanation-agnostic fine-tuning strategy to prevent information leakage and ii) a controlled random masking operation to overcome the OOD issues observed in the application of prior evaluation metrics. The fine-tuning process employs stochastic masking operations, such as randomly dropping pixels in images, tokens in language, or time steps in time series data, to generate augmented training samples. This augmented data is then used to fine-tune a surrogate model. Unlike previous approaches such as ROAR (Hooker et al., 2019), our strategy effectively mitigates the risk of information leakage and label bias by using explanation-agnostic stochastic masks, while also offering improved efficiency as it eliminates the need to retrain the model for each individual explainer. During the evaluation phase, F-Fidelity implements a removal strategy that generates stochastic masks conditioned on the explainer output, designed to be in-distribution with respect to the masks used in the fine-tuning step, which mitigates the OOD issues observed in the application of prior evaluation metrics.
@@ 2.3 Quantifying the Performance of Explainers
A key challenge in explainability research is to quantify and compare the performance of various explainers. The performance of an explainer can be formally quantified in terms the total variation distance $d_{TV}(P_{Y|\mathbf{X}},P_{Y|\mathbf{X}\odot\mathbf{M}})$ as a function of the average explanation size $\mathbb{E}_{\mathbf{X}}(\|\mathbf{M}\|_{1})$. However, in most problems of interest, the underlying statistics $P_{\mathbf{X},Y}$ is not available, and hence direct evaluation of the aforementioned total variation distance is not possible. As discussed in the introduction, some datasets are accompanied by ground-truth explanations, which enables the use of measures such as AUC and IoU for evaluating the quality explainers. However, the ground-truth explanations are available only for a limited collection of datasets, and even when ground-truth explanations are available, they may not accurately reflect the model’s internal decision-making processes. To address the aforementioned issues, a widely used set of metrics have been proposed in the literature which are based on the removal strategy. In CV, two removal orders have been considered (Samek et al., 2016; Yuan et al., 2022): MoRF, which evaluates explanations by removing the most influential pixels first, and LeRF, which begins with removing the least influential pixels. These approaches provide complementary perspectives on feature importance - MoRF assesses whether removing important features significantly impacts predictions, while LeRF verifies if retaining important features is sufficient for model performance. In the graph domain, an alternative but conceptually related removal strategy is used, where first the size of explanation subgraphs is determined according to a sparsity parameter, and then edges are removed either from the explanation subgraph of the desired size, or the non-explanation subgraph which is its complement. In this paper, we use metrics based on this graph domain removal strategy, called the Fidelity (Fid) metric in the literature (Pope et al., 2019; Bajaj et al., 2021; Yuan et al., 2022). The Fidelity metrics align with MoRF and LeRF principles through $Fid^{+}$ and $Fid^{-}$ respectively. Formally, given the input and label pair $(\mathbf{x},y)$ and a binary mask $\mathbf{m}$, the $Fidelity$ metrics are defined as follows:
$Fid^{+}(\psi)$ $=\frac{1}{|\mathcal{T}|}\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{T}}\mathds{1}(y_{i}=f(\mathbf{x}_{i}))-\mathds{1}(y_{i}=f(\mathbf{x}_{i}-\mathbf{x}_{i}\odot\mathbf{m}_{i}))$ (1)
$Fid^{-}(\psi)$ $=\frac{1}{|\mathcal{T}|}\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{T}}\mathds{1}(y_{i}=f(\mathbf{x}_{i}))-\mathds{1}(y_{i}=f(\mathbf{x}_{i}\odot\mathbf{m}_{i})),$ (2)
@@
where $\mathcal{T}$ is the dataset used for evaluating the performance of the explainer, $n$ is the size of the dataset, $\mathds{1}(\cdot)$ denotes the indicator function, and $\mathbf{m}_{i}=\psi(\mathbf{x}_{i})$ is the explanation corresponding to $\mathbf{x}_{i}$ produced by the explainer $\psi(\cdot)$. Here, $Fid^{+}$ measures prediction changes when removing important features (similar to MoRF), while $Fid^{-}$ evaluates model performance when keeping only important features (similar to LeRF).
A significant limitation of explanation evaluation through removal strategies is the OOD problem (Hooker et al., 2019; Zheng et al., 2024). When we remove elements from an input - whether they are pixels in images, time steps in time series, or edges in graphs - the modified input may no longer follow the original data distribution that the model was trained on. For instance, when evaluating image explanations by zeroing out important pixels, the resulting images with black patches are unlikely to resemble natural images. Consequently, the model’s predictions on these modified inputs may be unreliable, not due to low quality of the explanation itself, but because the model is operating outside its training distribution.
This OOD problem has been analyzed through different approaches. In ROAR (Hooker et al., 2019), which focused on CV tasks, the solution was to retrain the model on perturbed data where important input elements identified by the explainer are removed. While this approach mitigates the OOD issue, Rong et al. (2022) identified that it suffers from information leakage through the binary masks of removed pixels - the pattern of removals itself can contain enough class information to affect the evaluation outcome. Moreover, ROAR requires retraining a separate model for each explainer being evaluated, making it computationally expensive and time-consuming when comparing multiple explanation methods.
@@
Similarly, In R-Fidelity (Zheng et al., 2024), which considered the evaluation of GNN explainers, it was argued that the Fidelity metric highly relies on the robustness of the underlying classifier to removal of potentially large sections of the input, e.g.,the removal of a large subgraph explanation for $Fid^{+}$ or its complement for $Fid^{-}$. That is, the classifier should be robust to OOD inputs for the Fidelity metric to align with those of the (theoretically justified) total-variation-based metric discussed in the previous sections. Rather than retraining the model, R-Fidelity introduces a stochastic removal strategy that addresses the OOD issue by controlling the size of removed sections and randomly sampling which elements to remove, thus limiting the distribution shift of perturbed inputs. Specifically, the following Robust Fidelity metrics (RFid) was introduced:
$RFid^{+}(\psi,\alpha^{+},s)$ $=\frac{1}{|\mathcal{T}|}\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{T}}\mathds{1}(y_{i}=f(\mathbf{x}_{i}))-P(y_{i}=f(\chi^{+}(\mathbf{x}_{i},\alpha^{+},s))$ (3)
$RFid^{-}(\psi,\alpha^{-},s)$ $=\frac{1}{|\mathcal{T}|}\sum_{(\mathbf{x}_{i},y_{i})\in\mathcal{T}}\mathds{1}(y_{i}=f(\mathbf{x}_{i}))-P(y_{i}=f(\chi^{-}(\mathbf{x}_{i},\alpha^{-},s)),$ (4)
@@
where $\chi^{+}(\mathbf{x}_{i},\alpha^{+},s)$ is a sampling function which randomly, uniformly, and independently removes $\lfloor s\alpha^{+}\rfloor$ elements from the $s$ highest scoring elements of $\mathbf{x}_{i}$ based on the scores produced by $\phi(\mathbf{x}_{i})$, and $\chi^{-}(\mathbf{x}_{i},\alpha^{-})$ removes $\lceil(td-s)\alpha^{-}\rceil$ elements from the lowest scoring $td-s$ elements. If $\alpha^{+}=\alpha^{-}=1$, then the RFid metric reduces to the Fid metric. On the other hand, as $\alpha^{+}$ and $\alpha^{-}$ are decreased, fewer input elements are removed, hence requiring lower OOD robustness to ensure the accuracy of the evaluation output.
3 Robust Fidelity via fine-tuning and stochastic removal
@@
A partial solution in the graph domain addresses this issue by removing only an $\alpha^{+}$ fraction of the explanation subgraph and $\alpha^{-}$ fraction of the non-explanation subgraphs (Zheng et al., 2024). However, we argue that two issues degrade the evaluation quality of the RFid metric. First, the classifier may lack robustness and produce unreliable outputs even when the input is only slightly perturbed. Second, if the original explanation size is large (small), then removing an $\alpha^{+}$ ($\alpha^{-}$) of the explanation (non-explanation) part of the input, this would still yield OOD inputs.
In this work, we introduce a simple yet effective framework, F-Fidelity, for robust evaluation of XAI methods. Our strategy can be summarized as fine-tuning and stochastic removal. Specifically, we first fine-tune the model with randomly masked inputs to improve its robustness to perturbation. Then employ a controlled stochastic removal process that ensures the perturbed inputs remain within the distribution seen during fine-tuning.
@@
To achieve reliable predictions on partially removed inputs, we design a fine-tuning process that randomly removes up to $\beta\in[0,1]$ ratio of input elements. To elaborate, we introduce a stochastic mask generator $P_{\beta}:(t,d)\mapsto\mathbf{M}_{\beta}$, which takes the input dimensions $(t,d)$ as input and outputs a mask $\mathbf{M}_{\beta}\in\{0,1\}^{t\times d}$ of size $\beta td$, i.e. with up to $\beta{td}$ non-zero elements. For instance, in image classification, the mask generator is designed to select random image pixels or patches for removal. Formally, we define the fine-tuning loss as:
$L=\mathbb{E}_{\mathbf{X},Y}\left[\mathcal{L}\left(f\left(\mathbf{X}-P_{\beta}\odot\mathbf{X}\right),\ Y\right)\right],$ (5)
where $\mathcal{L}$ is the loss function used during training (e.g., cross-entropy).
In the evaluation process, we modify the RFid metric to ensure consistency with our fine-tuning strategy by upper-bounding the total number of removed elements by $\beta td$ — the same bound used during fine-tuning. That is, for a fixed $\beta$ , and RFid parameters $\alpha^{+}_{orig},\alpha^{-}_{orig}\in[0,1]$, we set the upper-bounded RFid parameters as
$\alpha^{+}=\min(\alpha^{+}_{orig},\frac{\beta td}{s})\quad\text{and}\quad\alpha^{-}=\min(\alpha^{-}_{orig},\frac{\beta td}{(td-s)}),$ (6)
so that the sampling functions $\chi^{+}$ and $\chi^{-}$ remove the minimum of $\alpha^{+}_{orig}s$ (based on explanation size) and $\beta td$ (based on input size) elements for $\chi^{+}$, and the minimum of $\alpha^{-}_{orig}(td-s)$ and $\beta td$ elements for $\chi^{-}$, providing absolute upper bounds on the number of removed elements.
@@
The pipeline of our method is shown in Algorithm 1. We denote the resulting metrics, which use the fine-tuning process and the $RFid^{+}$ and $RFid^{-}$ metrics with sampling rates that are truncated based on $\beta$ ( equation 6), as $FFid^{+}(\psi,\alpha^{+}_{orig},\beta,s)$ and $FFid^{-}(\psi,\alpha^{-}_{orig},\beta,s)$, respectively.
It is worth noting that both $FFid^{+}$ and $FFid^{-}$ can take negative values in certain cases. This occurs when the accuracy after masking exceeds the original prediction accuracy. Alternative formulations could enforce positive values by only reporting masked accuracy like ROAR (Hooker et al., 2019) and deletion/insertion scores (Petsiuk et al., 2018; Pan et al., 2021).
4 Determining the Explanation Size via Fidelity Metrics
A critical challenge in explainable AI is determining the appropriate size or scope of explanations. Ideally, ground truth explanations would be discretized into distinct clusters representing different levels of importance. For instance, in image classification, pixels associated with the target object tend to receive high importance scores. Conversely, pixels corresponding to the background or irrelevant regions receive low scores. However, in many practical scenarios, even good explainers that produce accurate explanation masks — as measured by the Fid and RFid evaluation metrics — may yield explanation scores that are not discretized into distinct clusters.
In this section, we theoretically demonstrate that our proposed evaluation metric can recover the cluster sizes given an explainer that outputs the correct explanation mask (i.e., correctly ranks the importance of input elements). Thus, provided the explainer outputs an accurate mask function, our metric can recover the explanation size (also known as sparsity). To provide a concrete theoretical analysis, we first consider a classification problem under a set of idealized assumptions that generalize the above observations on the clustering of explanation scores. Specifically, let us consider a classification task defined by a joint distribution $P_{\mathbf{X},Y}$ and a classifier $f:\mathbf{x}\mapsto y$. We assume that the input elements can be partitioned into several influence tiers. That is, for any given input $\mathbf{x}$, there exists a partition $\mathcal{C}_{k}(\mathbf{x}),k\in[r]$ of the index set $[t]\times[d]$, where $\mathcal{C}_{k}(\mathbf{x})$ represents the set of indices of the input elements belonging to tier $k$, and $c_{k}=|\mathcal{C}_{k}(\mathbf{x})|$ are the (fixed) tier sizes. For a given mask $\mathbf{m}$, the probability of correct classification based on the masked input $\mathbf{x}\odot\mathbf{m}$ depends only on the counts of unmasked elements in each influence tier. Formally, $P(Y|\mathbf{x}\odot\mathbf{m})=g(j_{1},j_{2},\dots,j_{r})$, where $g:[c_{1}]\times[c_{2}]\times\cdots\times[c_{r}]\to[0,1]$ is a function monotonically increasing with respect to the lexicographic ordering on its input, and $j_{k}\in[c_{k}]$ is the number of elements in $\mathcal{C}_{k}(\mathbf{x})$ whose corresponding mask element in $\mathbf{m}$ is non-zero (unmasked). We further focus on Shapley-value-based explanations, which provide a theoretical foundation for our analysis. Recall that, given label $y$, the Shapley value associated with an element $(i,j)\in[t]\times[d]$ of $\mathbf{x}$ is given as (Lundberg, 2017):
$S_{\mathbf{x}}(i,j)=\sum_{\mathbf{m}:m_{i,j}=0}\frac{\|m\|_{1}!(td-\|m\|_{1}-1)!}{(td)!}(P(Y=y|\mathbf{x}\odot\mathbf{m}^{\prime})-P(Y=y|\mathbf{x}\odot\mathbf{m})),$
Theorem 1.
The proof is provided in the Appendix.
This theorem shows that $FFid^{+}$ can recover the size of the most influential tier (i.e., the first cluster size) when the explainer’s ranking is close to that of an ideal Shapley-based explainer. Specifically, the value of $s$ in which $FFid^{+}$ changes direction corresponds to the size of an influence tier. This is verified in the empirical evaluation provided in Section D.8. This result implies that even when an explainer provides continuous scores without distinct clustering, our metric can infer the underlying discrete structure of the ground truth explanations and recover the explanation size.
5 Experiments